OpenStudy (anonymous):

The weak acid HQ has pKa of 4.89. Calculate the (OH-) of 0.500 M HQ

OpenStudy (aaronq):

First find \([H^+]\), then use \(\sf [H^+][OH^-]=K_w=1.0*10^{-14}\)

OpenStudy (cuanchi):

\[pKa = -\log Ka => Ka = 10^{-pKa}\] \[[H ^{+}]= \sqrt{Ka \times Ca}\] where Ca is the initial concentration of the acid assuming that the "x too small"